OLL and PLL skip chances

Lt-UnReaL

Veteran
Joined
Dec 6, 2006
Messages
1,146
Location
Rochester, NY
WCA ID
2008CLAY01
Apparently the chances for an OLL skip is 1/216 solves. I've never gotten an OLL skip in any speedsolve in the passed 5 months, and I solve about 30 times a day.

This one also comes up 1/216. I've gotten this OLL about 4 times in the last 3 days, about 90 solves.
oll03.gif

That's equivalent to 1/22.5 solves...Of course I was just lucky, but I have gotten this at least 10 times in the passed 5 months, and no OLL skips...

Now the chances for PLL skips are 1/72. I've gotten about 3-4 PLL skips in the passed 5 months, but at least 15-20 N perms(about 7-10 for each one). They also come up 1/72...

How can these numbers be right? How can people get OLL/PLL skips all the time and I never get them... :(
 
You do know that these probablities are not-mutual.
You could do 10,000 solves and get a LL skip everytime, it is unlikely but it could happen because one solve does not affect the probablities of the next solve.

The probability for an OLL skip on one solve is 1/216, you can't do 216 solves and expect an OLL skip. The probabilities come from the states of the cube, where 1 out of 216 states is an OLL skip.
 
wow...
I never got OLL skips either excpet when I do VHF2L but then my chances are a little different.
I get 2 or 3 PLL skips a day but I don't time most of my solves since they are just to keep me entertained as I walk 15 minutes to class. But I solve a lot more cubes a day than you.
 
The probability for an OLL skip on one solve is 1/216, you can't do 216 solves and expect an OLL skip.
Wrong! One OLL skip is *exactly* what you should expect.

I don't understand. What I use to understand is that 1/216 expressed the probabilities of an OLL skip but lets say the case 158 repeated a lot so I wouldn't have a OLL skip until the other 215 cases have been experienced.
 
yeah, Stefan is right. But I think Joey's point is that if you didn't have an OLL skip in solve 1, you shouldnt expect the chances for an OLL skip in solve 2 to be reduced. The probability is a constant 1/216

Dan
 
You have more frequent OLL skips if you can (partially) control LL edge orientation during F2L.
There are some regular VHF2L cases that you can learn to orient LL during F2L.
 
I've had 3 or 4 OLL skips in the one and a half months I've been cubing with fridich, I've had about 5 PLL skips, they always happen when I'm not timing myself, just when I'm messing around.
 
yea in my last 5 solves, i hade 2 pll skips, and one oll skip, well i did one lol i think, cant remember, but quite odd. nice times for a noob tho:D
 
The probability for an OLL skip on one solve is 1/216, you can't do 216 solves and expect an OLL skip.
Wrong! One OLL skip is *exactly* what you should expect.

I suppose it was the way I worded it.

You should *expect* one OLL skip, but it does not mean you *will* get that one OLL skip.


Stefan: Correct?

Yeah, I knew this. I'm just saying the chances for a person to solve it like 1,500 times and only receive 1 OLL skip out of them all are pretty bad.

As the number of cubes you solve approaches infinity the ratio of OLL skips you get will approach 1/216
Ah, this makes more sense.
 
I'm pretty sure the odds of encountering either skip a single time can be represented by exponential decay functions. O(X) for OLL cases, P(X) for PLL cases, X for the number of solves.

O(X) = 1 - (215/216)^X
P(X) = 1 - (71/72)^X
 
i get way more PLL skips than OLL skips, excepted for the controlled ones.
 
Back
Top